OpenStudy (anonymous):

6. a^2b^6 2 ____________ * __________ = 4 a^2b^3

OpenStudy (skullpatrol):

$$\frac{a^2b^6}{4}*\frac{2}{a^2b^3}$$

OpenStudy (goformit100):

Sir Question is unclear .

OpenStudy (anonymous):

yes that is that is it

OpenStudy (skullpatrol):

The rule for simplifying fractions: If b,c and d are real numbers with b=/=0 and d=/=0, then $$\frac{bc}{bd}=\frac{c}{d}$$This rule allows you to divide the numerator and denominator of a fraction by the same nonzero number.

OpenStudy (anonymous):

so do i divide the top and bottom

OpenStudy (skullpatrol):

Sort of... first divide the numerator and the denominator by 2

OpenStudy (jack1):

when you're multiplying fractions, you times straight across (numerator x numerator) / denominator x denominator from there you get a single fraction once you have that fraction, you can cancel factors

OpenStudy (jack1):

(a^2 x b^6 x 2) / (4 x a^2 x b^3)

OpenStudy (anonymous):

ok I have that

OpenStudy (jack1):

2/4 can be reduced to 1/2 a^2 / a^2 can be cancelled to 1/1, which is just 1

OpenStudy (jack1):

and b^6 / b^3 can be reduced as 1/b^3 is the same as writing b^-3

OpenStudy (jack1):

so b^6 x b^-3 = b^(6-3) =b^3 following ok so far?

OpenStudy (anonymous):

yes i am following

OpenStudy (jack1):

cools, so with those cancelled, what are you left with?

OpenStudy (anonymous):

b^3

OpenStudy (jack1):

yeah, but remember the 2/4 reduced to 1/2... so you'll be left with b^3 / 2

OpenStudy (anonymous):

thank you so much

OpenStudy (jack1):

all good man, best of luck